# Bootstrap Off-policy with World Model

### v2

@@ 7 Acknowledgment
This research was generously supported by the Tsinghua University-Toyota Joint Research Center for AI Technology of Automated Vehicle. The core of this work was completed during the visit to the Mechanical System Control (MSC) Lab at the University of California, Berkeley. We extend our sincere gratitude to all the lab members for their constructive feedback during this period.
@@ B.2 Reproducibility Statement & Detailed Hyperparameters
In this paper, we evaluate each algorithm for each tasks over three random seeds. The CPU used is the AMD Ryzen Threadripper 3960X 24-Core Processor, and the GPU used is NVIDIA GeForce RTX 3090Ti. Taking Dog-run task in the DMC Suite as an example, the time taken to train 4M environment steps is around 50 hours.
For SAC and DreamerV3, we report the baseline results released by TD-MPC2, which are obtained from the official repositories of https://github.com/denisyarats/pytorch_sac and https://github.com/danijar/dreamerv3, respectively. For BMPC, we use the official implementation https://github.com/wertyuilife2/bmpc and align the settings such as environment interaction steps and evaluation protocol for a fair comparison.
@@
Number of value bins 101 Value dropout rate $1\%$
Warmup steps 1000
@@ C.2 Visualizations
C.3 Training Curves With 10M Environment Steps
We observe that the first two baselines perform very poor with 4M environment steps. Recall that they are: (1) SAC [haarnoja2018soft]: the state-of-the-art model-free off-policy RL algorithm under the maximum entropy framework; (2) DreamerV3 [hafner2025mastering]: the state-of-the-art imagination-driven model-based RL algorithm that learns from imaginary rollouts generated by a world model.
To enable a comprehensive comparison, we further evaluate SAC and DreamerV3 under an increased interaction budget of 10 million environment steps. The updated training curves are shown in Figure 15. Despite benefiting from up to five times more environment interactions, SAC still performs poorly, with TAR remaining close to zero on most tasks. DreamerV3 gradually surpasses TD-MPC2 and BMPC on certain tasks. Overall, our BOOM consistently outperforms all baselines across all evaluated tasks.
(a) AVG. DMC Suite
(b) Humanoid-stand
(c) Humanoid-walk
(d) Humanoid-run
(e) Dog-stand
(f) Dog-walk
(g) Dog-trot
(h) Dog-run
(i) AVG. Human. Bench
(j) H1hand-stand
(k) H1hand-walk
(l) H1hand-run
(m) H1hand-sit
(n) H1hand-slide
(o) H1hand-pole
(p) H1hand-hurdle

### v3

@@ 7 Acknowledgment
This study is supported by the Tsinghua University-Toyota Joint Research Center for AI Technology of Automated Vehicle, Beijing Natural Science Foundation (L257002), and SunRisingAI Lab.
@@ B.2 Reproducibility Statement & Detailed Hyperparameters
In this paper, we evaluate each algorithm for each tasks over three random seeds. The CPU used is the AMD Ryzen Threadripper 3960X 24-Core Processor, and the GPU used is NVIDIA GeForce RTX 3090Ti. Taking Dog-run task in the DMC Suite as an example, the time taken to train 2M iterations is around 50 hours.
For SAC and DreamerV3, we report the baseline results released by TD-MPC2, which are obtained from the official repositories of https://github.com/denisyarats/pytorch_sac and https://github.com/danijar/dreamerv3, respectively. For BMPC, we use the official implementation https://github.com/wertyuilife2/bmpc and align the settings such as number of iterations and evaluation protocol for a fair comparison.
@@
Number of value bins 101 Warmup steps 5000